I have a word problem....The product of three and a number decreased by seven is seventeen. What is the number? I need to formulate an equation. I wish word problem made since to me. Thank you
Answer by stanbon(75887)

I have a word problem....The product of three and a number decreased by seven is seventeen. What is the number?
------------------
EQUATION:
3(x-7) = 17
3x-21 = 17
3x = 38
x = 38/3
================================
Cheers,
Stan H.